# SI_appendix_phys

### v1

@@ Methods
    $d_{\text{H}}(t^{*})\geq 0.05$, $\tilde{c}(t^{*})\geq 0.01$ and $\sup_{t\in[t^{*}-500,t^{*}]}|\tilde{c}^{\prime}(t)|\geq 10^{-5}$to CP solution.
Traveling profiles: The profiles $\varrho_{\sigma}(z)$ and speed $c$ in Fig. 5: are obtained numerically by discretizing Eqs. (7,8) with second-order centered differences and solving the resulting zeroth-finding problem subject to mass constraints $\int\varrho_{0}\mathrm{d}z=\phi_{p}$ and $\int\varrho_{+}\mathrm{d}z=\int\varrho_{-}\mathrm{d}z=\phi_{a}/2$ in $[-L/2,L/2]$. In the finite-$L$ case ((7)), the system is initialized with a preexisting T solution with similar parameters or a steady state of (1) and solved subject to periodic boundary conditions and $\varrho(0)=\phi$ without loss of generality (since the problem has translational symmetry). In the large-domains case ($L\gg 1$), (8), the numerical procedure is as follows. Given parameter values $\phi$ and $\nu$ ((5)), we determine the inner solution via the coexisting phases procedure as above, resulting in liquid $(\varrho_{\sigma})_{l}$ and vapor $(\varrho_{\sigma})_{v}$ values the outer solution must match with (if $\phi_{v}(\nu)=\phi_{l}(\nu)$, it means that there is no inner region and we may proceed to solve (8) subject to periodic boundary conditions as in the finite-system case). Then (8) is discretized using second-order finite-differences and solved in $[0,L]$ subject to Dirichlet boundary conditions $\varrho_{\sigma}(0)=(\varrho_{\sigma})_{v}$ and $\varrho_{\sigma}(L)=(\varrho_{\sigma})_{l}$ and mass constraints as above. If no solution is found, it indicates there may be a second interface. We initialize with a previous single interface T solution with similar parameters. We then place a new interface at $z=\text{argmax}_{z}|\partial_{z}\varrho|$. In both finite $L$ and $L\gg 1$ cases, we solve the resulting systems of equations using the NonlinearSolve.jl package in Julia (see Appendix 2).
Acknowledgements.
We thank Mike Cates, Clement Erignoux, Tal Agranov, and Sarah Loos for helpful discussions. M. Bruna was supported by a Royal Society University Research Fellowship (grant no. URF/R1/180040). J. Mason was supported by the Royal Society Award (RGF/EA/181043) and the Cantab Capital Institute for the Mathematics of Information of the University of Cambridge.
@@ Acknowledgements.
A Hydrodynamic limit
The hydrodynamic system of PDEs (1) describes the local density of each type of particle as the mesh parameter $h\to 0$. For ease of reference, we rewrite (1) below in terms of $\rho=\rho_{+}+\rho_{-}+\rho_{0}$, $\rho_{a}=\rho_{+}+\rho_{-}$ and ${m}=\rho_{+}-\rho_{-}$:
$\partial_{t}\rho$ $=\nabla\cdot\nabla\rho-\text{Pe}\partial_{x}\left[(1-\rho){m}\right],$ (S1)
$\partial_{t}\rho_{a}$ $=\nabla\cdot\left[d_{s}(\rho)\nabla\rho_{a}+\rho_{a}{\cal D}(\rho)\nabla\rho\right]-\text{Pe}\partial_{x}\left[\rho_{a}s(\rho){m}+d_{s}(\rho){m}\right],$ (S2)
$\partial_{t}{m}$ $=\nabla\cdot\left[d_{s}(\rho)\nabla{m}+{m}{\cal D}(\rho)\nabla\rho\right]-\text{Pe}\partial_{x}\left[s(\rho){m}^{2}+d_{s}(\rho)\rho_{a}\right]-2{m}\;.$ (S3)
We define $\eta_{\sigma}(\bm{x},t)$ to be $1$ if there is a $\sigma$ particle at position $\bm{x}$ and 00 otherwise. The local density is formally defined as the mean number of particles in a mesoscopic box of radius $r$ around $\bm{x}$,
@@ A Hydrodynamic limit
for $1\gg r\gg h$. The existence of a hydrodynamic limit means that the random variables $\hat{\rho}_{\sigma}$ converge (in probability) to deterministic densities $\rho_{\sigma}$, which are solutions to the hydrodynamic PDE system Eqs. (S1-S3).
We obtain the hydrodynamic limit (S1) of the APLG by generalizing the work of Erignoux [21]. That work considered a system of pure active particles with continuously varying orientations $(\cos\theta,\sin\theta)$ with $\theta\in[0,2\pi)$ instead of only $\theta=\{0,\pi\}$ used here. In both cases, the proof of the convergence $\hat{\rho}_{\sigma}\to\rho_{\sigma}$ is technically challenging because the models are of non-gradient type in the sense of [20]. It is worth emphasizing that this classification is separate from whether the model can be derived as a gradient flow of an equilibrium free energy. It means instead that the current $j_{\bm{x},\bm{x}+h\bm{e}_{i}}$ between two sites $\bm{x}$ and $\bm{x}+h\bm{e}_{i}$, cannot be written as the discrete difference of a local function $g$
@@
The non-gradient method [24] involves projecting the current onto a space of discrete differences and proving that it can be replaced by its local average in the hydrodynamic limit, e.g.,
$j_{\bm{x},\bm{x}+h{\bm{e}}_{1}}\simeq d_{s}(\rho)[\eta_{\sigma}(\bm{x}+h\bm{e}_{1})-\eta_{\sigma}(\bm{x})]+{\cal D}(\rho)[\eta(\bm{x}+h\bm{e}_{1})-\eta(\bm{x})]+\sigma d_{s}(\rho)\eta_{\sigma}(\bm{x})+s(\rho)m\eta(\bm{x}),$ (S6)

### v2

@@ Methods
    $d_{\text{H}}(t^{*})\geq 0.05$, $\tilde{c}(t^{*})\geq 0.01$ and $\sup_{t\in\mathcal{T}^{*}}|\tilde{c}^{\prime}(t)|\geq 10^{-5}$to CP solution.
$\bullet$ Traveling profiles: The profiles $\varrho_{\sigma}(z)$ and speed $c$ in Fig. 5 are obtained numerically by discretizing Eqs. (7,8) with second-order centered differences and solving the resulting zeroth-finding problem subject to mass constraints $\int\varrho_{0}\mathrm{d}z=\phi_{p}$ and $\int\varrho_{+}\mathrm{d}z=\int\varrho_{-}\mathrm{d}z=\phi_{a}/2$ in $[-L/2,L/2]$. The finite-$L$ system (7) is initialized with a preexisting T solution with similar parameters or a steady state of (1) and solved subject to periodic boundary conditions and $\varrho(0)=\phi$ without loss of generality (since the problem has translational symmetry). For $L\gg 1$, the numerical procedure to solve (8) is as follows. Given parameter values $\phi$ and $\nu$ (5), we determine the inner solution via the abovementioned Coexisting Phases procedure. This results in liquid $(\varrho_{\sigma})_{l}$ and vapor $(\varrho_{\sigma})_{v}$ values, which become boundary values for the outer problem (8). [If $\phi_{v}(\nu)=\phi_{l}(\nu)$, it means that there is no inner region and we may proceed to solve (8) subject to periodic boundary conditions as in the finite $L$ case.] Then (8) is discretized using second-order finite-differences and solved in $[0,L]$ subject to Dirichlet boundary conditions $\varrho_{\sigma}(0)=(\varrho_{\sigma})_{v}$ and $\varrho_{\sigma}(L)=(\varrho_{\sigma})_{l}$ and mass constraints as above. If no solution is found, it indicates there may be a second interface. We initialize with a previous single interface T solution with similar parameters. We then place a new interface at $z=\text{argmax}_{z}|\partial_{z}\varrho|$. In both finite $L$ and $L\gg 1$ cases, we solve the resulting systems of equations using the NonlinearSolve.jl package in Julia (see Appendix 2).
Data availability
The particle and PDE simulations data that support the findings of this study have been deposited in the figshare repository [58].
Code availability
The code that supports the findings within this paper is available at the GitHub repository [59].
Acknowledgements
We thank Tal Agranov, Martin Burger, Mike Cates, Clement Erignoux, Sarah Loos, and Johannes Zimmer for helpful discussions. M. Bruna was supported by a Royal Society University Research Fellowship (grant no. URF/R1/180040). J. Mason was supported by the Royal Society Award (RGF/EA/181043) and the Cantab Capital Institute for the Mathematics of Information of the University of Cambridge. For the numerical work, we used the Julia programming language [60] and the following packages and tools: DrWatson.jl [61], DifferentialEquations.jl [62], NonlinearSolve.jl [63]. For the purpose of open access, the authors have applied a CC BY public copyright licence to any author accepted manuscript arising from this submission.
@@ Acknowledgements
Appendix A Hydrodynamic limit
The hydrodynamic system of PDEs (1) describes the local density of each type of particle as the lattice spacing $h\to 0$. For ease of reference, we rewrite (1) below in terms of $\rho=\rho_{+}+\rho_{-}+\rho_{0}$, $\rho_{a}=\rho_{+}+\rho_{-}$ and ${m}=\rho_{+}-\rho_{-}$:
$\partial_{t}\rho$ $=\nabla\cdot\nabla\rho-\text{Pe}\partial_{x}\left[(1-\rho){m}\right],$ (S1a)
$\partial_{t}\rho_{a}$ $=\nabla\cdot\left[d_{s}(\rho)\nabla\rho_{a}+\rho_{a}{\cal D}(\rho)\nabla\rho\right]-\text{Pe}\partial_{x}\left[\rho_{a}s(\rho){m}+d_{s}(\rho){m}\right],$ (S1b)
$\partial_{t}{m}$ $=\nabla\cdot\left[d_{s}(\rho)\nabla{m}+{m}{\cal D}(\rho)\nabla\rho\right]-\text{Pe}\partial_{x}\left[s(\rho){m}^{2}+d_{s}(\rho)\rho_{a}\right]-2{m}.$ (S1c)
To obtain these equations from the particle-based dynamics of the APLG, we define $\eta_{\sigma}(\bm{x},t)$ to be $1$ if there is a $\sigma$-particle at position $\bm{x}$ and 00 otherwise. The local density is formally defined as the mean number of particles in a mesoscopic box of radius $r$ around $\bm{x}$,
@@ Appendix A Hydrodynamic limit
for $1\gg r\gg h$. Recalling that the lattice is embedded in a physical domain of size $\ell_{x}\times\ell_{y}$ and that the total numbers of active (and passive) particles per site are $\phi_{a}$ (and $\phi_{p}$), the relevant limit is $\ell_{y}\to 0$ at fixed $\ell_{x},\ell_{y},\phi_{a},\phi_{p}$. The hydrodynamic limit exists if that the random variables $\hat{\rho}_{\sigma}$ converge (in probability) to deterministic densities $\rho_{\sigma}$, which are solutions to the hydrodynamic PDE system (S1).
This is proven in the APLG by generalizing the work of Erignoux [19]. That work considered a system of pure active particles with continuously varying orientations $(\cos\theta,\sin\theta)$ with $\theta\in[0,2\pi)$ instead of only $\theta=\{0,\pi\}$ used here. In both cases, the proof of the convergence $\hat{\rho}_{\sigma}\to\rho_{\sigma}$ is technically challenging because the models are of nongradient type in the sense of [18]. It is worth emphasizing that this classification is separate from whether the model can be derived as a gradient flow of an equilibrium free energy. Indeed, the hydrodynamic limit of a mixture of two species undergoing a symmetric simple exclusion process is a gradient flow, but the underlying microscopic model is of nongradient type [26]. Being of nongradient type means instead that the current $j_{\bm{x},\bm{x}+h\bm{e}_{i}}$ from site $\bm{x}$ to $\bm{x}+h\bm{e}_{i}$, cannot be written as the discrete difference of a local function $g$
@@
The original proof in [19] is extremely technical and long; for an abridged version, we refer the reader to [25, Sec. 5].
@@
The nongradient method [26] involves projecting the current onto a space of discrete differences and proving that it can be replaced by its local average in the hydrodynamic limit, e.g., the instantaneous current of $\sigma$-particles going between neighboring sites $\bm{x}$ and $\bm{x}+h\bm{e}_{1}$ can be approximated as
$j_{\bm{x},\bm{x}+h{\bm{e}}_{1}}^{\sigma}\simeq d_{s}(\rho)[\eta_{\sigma}(\bm{x}+h\bm{e}_{1})-\eta_{\sigma}(\bm{x})]+{\cal D}(\rho)[\eta(\bm{x}+h\bm{e}_{1})-\eta(\bm{x})]+\text{Pe}\left[\sigma d_{s}(\rho)\eta_{\sigma}(\bm{x})+s(\rho)m\eta_{\sigma}(\bm{x})\right],$ (S4)